OpenStudy (anonymous):

x-2/x-4 = x-6/x-9 solve for x .

OpenStudy (anonymous):

cross multiply this u will get the x

OpenStudy (anonymous):

duuddeee , its soo hard . i get confused .

OpenStudy (anonymous):

ok let me do it

OpenStudy (anonymous):

x^2-2x-9x+18=x^2-4x-6x+24 2x-3x-6=0 -x=6 x=-6